OpenStudy (anonymous):

f(x) kcosx/pi-2x, if x=/= pi/2 3, if x = pi/2 Function f is continuous at x = pi/2 :( I can't do this type of continuity problem :( @rational

OpenStudy (anonymous):

Hi rational :) I'm brushing up everything today :)

OpenStudy (anonymous):

The reason I got stuck in this one is when it is not equal to pi/2, what should I substitute x as?

OpenStudy (rational):

heyy its no different from previous problems : LHL = RHL = f(pi/2)

OpenStudy (anonymous):

What do I substitute x as? when it is not equal to pi/2?

OpenStudy (rational):

find the limit as x->pi/2

OpenStudy (anonymous):

but shouldn't you not use pi/2? its given in that right? unfortunately, this is one of the problems they asked in the exam last time.. I'm sure this exact type will come back this time

OpenStudy (rational):

you're given `f(pi/2) = 3` and everywhere else the function is given by `f(x) = kcosx/pi-2x` for the function to be continuous at x=pi/2, the limit must equal the function value

OpenStudy (anonymous):

so your supposed to plug in pi/2 regardless?

OpenStudy (rational):

for the function to be continuous at \(x=\pi/2\), we must have \[\lim\limits_{x\to \pi/2}~f(x) = f(\pi/2)\]

OpenStudy (anonymous):

so if we write kcos(pi/2) directly, it will give a 0 since cos(pi/2) = 0 and the whole thing will be zero..

OpenStudy (anonymous):

oh we're taking a little less than pi/2

OpenStudy (rational):

whats exactly the expression ? can you use latex/parenthesis..

OpenStudy (anonymous):